nework Assignment Score: 55.00 Submit Assignment for Gradin Question 8 of 10 Check My Work Chapter 10 Homework Save Problem 10.13 10-6: Cost of New Common Stock, e Cost of Common Equity with Flotation Ballack Co.s common stock currently sells for $4 expected long-run dividend payout ratio is 20%, and the expected return on equity (ROE is 12%. New stock can be sol flotation cost of 15% would be incurred, what would be the cost of new equity? Round your answer to two decimal places. 8.75 per share. The growth rate is a constant 9.6%, and the company has an expected dividend yi d to the public at the current price, but a 03 9.64 Hide Feedback Incorreet
Answer #1

Maximum sustainable growth rate (g) = 9.6 %

This can be confirmed based on
Growth Rate = ROE x (1 - payout ratio)
Growth Rate = 0.12 x (1-0.20)
Growth Rate = 0.096 or 9.6%

Common Stock Price = 48.75
Flotation cost = 15% of 48.75 = 7.3125
Net Proceed = 48.75 - 7.3125 = 41.4375

Dividend Yield = 4%
D1/Po = 4%
D1/48.75 = 4%
D1 = 4%*48.75 = 1.95

Now, As Dividend Growth Model

Cost of New Equity = D1/Net Proceed + Growth Rate
Cost of New Equity = 1.95/41.4375 + 9.6%
Cost of New Equity = 4.71% + 9.6%
Cost of New Equity = 14.31%